Questions on latent heat

Specific latent heat of fusion of water = 334 000J/kg
Specific latent heat of vaporisation of water = 2 265 000J/kg
Specific heat capacity of water = 4200J/kg
E = mL
E = m x c x 
You might need either of these equations to solve the answers.


1) Calculate the mass of water that turns into a gas when heated. The water receives 895kJ of thermal energy.

2) How much energy does it take to melt 0.2kg of ice?

3) Calculate the specific latent heat of fusion for 0.6kg of a substance that melts when 45 000J of thermal energy is supplied to it.

4) How much energy is needed to vaporise 0.3kg of water?


5) How much energy would be needed to melt 20g of ice?
1
Expert's answer
2019-10-25T11:30:24-0400

1)


"m=\frac{Q}{L}=\frac{895000}{2265000}=0.395\ kg"

2)


"Q=mL=(0.2)(334000)=66800\ J=66.8\ kJ"

3)


"L=\frac{Q}{m}=\frac{45000}{0.6}=75000\frac{J}{kg}"

4)


"Q=mL=(0.3)(2265000)=679500\ J"

5)


"Q=mL=(0.02)(334000)=6680\ J"